Q No 1) The Weaving Department of a factory occupies 3.5 times as much floor as the Shipping Department. If the Weaving Department occupies 4,200 square feet, what is space occupies by the Shipping Department?
Q No 2) A manufacturer of a certain product knows that he can sell as many units as he produces in a week at a price of Rs. 10 each. His variable cost of production is Rs. 2 per unit and total fixed cost for a week is Rs. 8,000. Determine break-even point graphically in units and value.

Answer by ikleyn(52787)

.

Q No 2.
----------

The basic equation is


    10x = 8000 + 2x


The solution is

    10x - 2x = 8000

     8x      = 8000

      x      = 8000/8 = 1000.


ANSWER.  Break-even point is 1000 units.